# On the Resilience of Multi-Agent Systems with Malicious Agents

### v2

@@ 1 Introduction
Multi-agent collaboration has further boosted Large Language Models’ (LLMs) already impressive performance across various downstream tasks, including code generation Liu et al. (2023); Lee et al. (2024), math problem solving Lu et al. (2024); Liang et al. (2024), and text translation Jiao et al. (2023); Wu et al. (2024). In such multi-agent systems, improvements are achieved by decomposing complex tasks into smaller, specialized sub-tasks handled by expert agents in a role-specific manner Chen et al. (2024); Li et al. (2024).
@@
However, the decentralized nature of multi-agent systems leaves them vulnerable to faulty or malicious agents, which could undermine or destroy collaboration. Consider a scenario where companies specializing in different areas produce expert agents, the lack of centralized control means that the multi-agent system may contain agents from various sources, some of which could be faulty or malicious. In a multi-agent coding system like Camel Li et al. (2023), a malicious coding agent could produce buggy code, causing severe errors or harmful outputs when executed by another agent.
@@
This paper studies the resilience of multi-agent systems against malicious agents, specifically their ability to recover from errors. Our focus extends beyond the transformation of agents into malicious ones for various tasks to their macro-level impact on collaborative dynamics, particularly how their presence leads to an overall performance decline in different systems. Recent studies Zhang et al. (2024); Tian et al. (2023); Amayuelas et al. (2024); Ju et al. (2024) have increasingly focused on safety issues within multi-agent systems. However, these studies mainly investigate attacks on agents to induce toxicity in their outputs or misinformation spread among all agents. While they assess malicious agent behavior against safety benchmarks like AdvBench Zou et al. (2023), they overlook the disruption of collaboration in solving general tasks and the impact of varying system structures.
Figure 1: What is the resilience of different systems against malicious agents on various tasks?
@@
To simulate the behaviors of malicious agents to study system resilience, we devise two approaches to corrupting benign ones, namely AutoTransform and AutoInject. AutoTransform transforms a given agent’s profile into a malicious version that retains original functionalities while introducing stealthy errors. AutoInject is designed to directly and automatically inject errors into messages spread among agents. The two methods offer automate introduction of errors in multi-agent systems without requiring manual modifications.
@@
To study the impact of different structures on system resilience, we select six multi-agent collaboration systems representing three classical real-world structures: Linear Hong et al. (2024); Dong et al. (2024), Flat Li et al. (2023); Wang et al. (2024c), and Hierarchical Chen et al. (2024); Liang et al. (2024). We evaluate the performance of these systems across four tasks: code generation Chen et al. (2021), math problem solving Liang et al. (2024), translation He et al. (2020), and text evaluation Wang et al. (2024a), as shown in Fig. 1. In our results, we find that the ierarchical structure exhibits the least performance degradation at $23.6\%$, aligning with its prevalence in human societal organizational structures Mihm et al. (2010). Code generation, as a relatively objective task, is most affected by malicious agents, experiencing a performance drop of $39.6\%$. Additionally, increasing the ratio of erroneous messages and using semantic errors results in a greater performance drop than increasing the number of errors per message and using syntactic errors. We also analyze the impact of roles and rounds on system resilience, which appears minimal.
@@
As an initial investigation into approaches for enhancing system resilience and defending against malicious agents, we introduce two strategies, each related to a specific error introduction method. The “Challenger” method adds to each agent’s profile the ability to challenge received messages, mirroring AutoTransform which rewrites agents’ profiles to make them malicious. The “Inspector” agent reviews and corrects messages, mirroring AutoInject which intercepts and injects errors into messages. We apply these defense methods to the two weaker systems: Self-collab with a linear structure and Camel with a flat structure. Our results demonstrate that both methods enhance system resilience, recovering up to $87.9\%$ of performance lost due to malicious agents.
The contribution of this paper can be summarized as follows:
@@
  • We explore the under-explored scenario where malicious agents exist and disrupt the collaboration, and are the first to examine how different structures of multi-agent systems affect resilience.
  • We implement AutoTransform and AutoInject to automatically introduce malicious agents, and design defense methods, the Inspector and the Challenger.
@@
  • We conduct extensive experiments involving six multi-agent systems across three system structures, applied to four common downstream tasks. Various factors that may influence resilience are analyzed, offering detailed insights into designing resilient multi-agent systems.
@@ 2 Preliminaries
A Management Science Perspective on Multi-Agent Systems
@@
In human collaboration, the capacity to mitigate mistakes or intentional disruptions within a team and maintain functionality despite individual failures is usually referred to as “resilience” Alliger et al. (2015); Boin & Van Eeten (2013); Hartwig et al. (2020). Resilience reflects the ability to handle internal errors, maintaining overall operation without being affected by a single failure. LLM-based multi-agent systems face safety issues where malicious agents produce errors too stealthy to be found by other agents but can cause undesired consequences. Therefore, holding this same ability as human collaboration becomes critical.
@@ 3 Methodology: Introducing Errors
Figure 2: Overview of our error introducing process. (a) Task information. (b) Multi-agent collaboration system without malicious agents. (c) AutoTransform modifies agent’s profile to turn it into malicious with preserving original functionalities. (d) AutoInject intercepts messages between agents and adds errors into the messages.
We offer two methods for introducing errors in multi-agent systems: AutoTransform converts agents into malicious entities that generate errors autonomously, while AutoInject directly introduces errors into messages. In this section, we first discuss the rationale behind designing the autonomous transformation agent in §3.1. Next, we introduce the method for directly injecting errors into messages within multi-agent systems in §3.2. These two methods are designed to be general-purpose, applicable to any agent profiles and downstream tasks. For presentation clarity, we use “message” to refer to intermediate outputs between agents, and “result” to denote the final output from the last agent.
3.1 AutoTransform: Malicious Agent Transformation
AutoTransform is an LLM-based approach that takes any agent’s profile as input and outputs a profile of a malicious agent performing the same functions but introducing stealthy errors. Drawing inspiration from how we manually convert an agent into malicious one, the design of AutoTransform follows three key steps: (1) To ensure applicability to any target agent and downstream tasks, AutoTransform first analyzes the input agent profile and extract the assigned task. This step helps to extract the task and identify potential ways to produce erroneous outputs. (2) Based on the task analysis, AutoTransform lists all possible methods to inject errors, emphasizing the need for stealth to avoid detection by other agents. (3) AutoTransform then rewrites the agent’s profile with these error-injection methods, ensuring that the original functionalities of the agent remain unchanged. An example of using AutoTransform to alter an agent’s profile is shown in Fig. 2c. The complete prompt is provided in §B.3 in the appendix.
@@ 3.2 AutoInject: Direct Error Injection
While AutoTransform can conveniently generate malicious agents, it is hard to ensure these agents introduce a specific number and type of errors due to the inherent randomness of the generation process. For example, “injecting syntax errors in 20% lines of the generated code” cannot be guaranteed by the malicious agents. However, precise error generation is crucial for analyzing the impact of various factors on system resilience. To address this, we introduce AutoInject, an approach that takes the outputs of other agents and intentionally injects specific errors. This approach allows for exact control over the proportion of erroneous messages, the specific errors within a message, and the types of errors introduced. We start by discussing two key factors in our study: error rate and error type.
@@
In this paper, we examine two aspects of error injection in multi-agent collaboration systems: Macro Perspective: We control the ratio of erroneous messages produced by a malicious agent in all its messages, which is a practical way to obscure its identity while facilitating stealthy errors. We denote this probability that a message is intentionally flawed as $P_{m}$. Micro Perspective: We manage the degree of error within each faulty message. For instance, in code generation tasks, we can adjust the number of errors per line of code. The proportion of a message that is erroneous is denoted by $P_{e}$.
@@
AutoInject is designed for specific tasks and agents. First, we assign the task, agent, error rates ($P_{m}$ and $P_{e}$), and error type. AutoInject then selects messages [^1] The final result message is excluded to allow system recovery. from the agent with a probability of $P_{m}$ and injects errors into $P_{e}$ of the total lines or sentences in the selected message. Errors are introduced automatically using LLMs, which receive the task introduction, error type, and the specific line or sentence. The LLMs produce erroneous lines or sentences, which replace the originals. An example of using AutoInject to modify an agent’s output into erroneous is shown in Fig. 2d. Prompts for different tasks are detailed in §B.4 the appendix.
@@ 4 Experiments
This section focuses on answering the following Research Questions (RQs):
  1. RQ1.
    Which of the three multi-agent system architectures exhibits the highest resilience (§4.2)?
  2. RQ2.
    Do different downstream tasks vary in their resilience to errors (§4.3)?
  3. RQ3.
    How do varying error rates (both $P_{m}$ and $P_{e}$) impact system resilience (§4.4)?
  4. RQ4.
    How do the two types of errors influence system resilience (§4.5)?
4.1 Settings
@@
We evaluate general-purpose task-solving abilities using common tasks:
  • Code Generation: HumanEval Chen et al. (2021) contains 164 hand-written programming problems to assess LLMs’ ability to synthesize correct and functional Python code. Accuracy (Pass@1) is used for evaluation.
  • Math Problem Solving: CIAR Liang et al. (2024) presents 50 questions with hidden traps to evaluate LLMs’ Counter-Intuitive Arithmetic Reasoning abilities, requiring multi-step reasoning. Accuracy is used for evaluation.
  • Translation: CommonMT He et al. (2020) consists of paired sentences to test models’ handling of three types of commonsense reasoning, especially in ambiguous contexts. We randomly sampled 100 sentences from the most challenging type, Lexical, for our evaluation, using BLEURT-20 Sellam et al. (2020); Pu et al. (2021) for evaluation, following the practice in Liang et al. (2024).
  • Text Evaluation: FairEval Wang et al. (2024a) includes 80 human-annotated “win/tie/lose” outcomes comparing responses from ChatGPT and Vicuna-13B, aiming to determine if the model’s preferences align with human judgments. Accuracy is used for evaluation.
@@
We consider three types of system architectures mentioned in §2:
  • Linear: MetaGPT Hong et al. (2024) employs Standard Operating Procedures (SOPs) to create an efficient workflow in a software company setting, utilizing five agents for code generation. Self-collaboration Dong et al. (2024) designs three roles, namely analyzers, coders, and testers, implemented using 2-5 agents on code generation task.
  • Flat: Camel Li et al. (2023) presents a framework where a “User” agent iteratively refines outputs from an “Assistant” agent, applicable across various tasks. SPP Wang et al. (2024c) uses Solo-Performance-Prompting to engage a single model into three personas for coding tasks.
  • Hierarchical: MAD Liang et al. (2024) introduces a Multi-Agent Debate framework with two debaters and one judge to promote divergent thinking in LLMs for various tasks. AgentVerse Chen et al. (2024) employs a dynamic recruitment process, selecting agents for multi-round collaboration as needed, utilizing four agents for our selected tasks.
Not all systems are designed to support the four tasks studied in this paper. Therefore, we modified the prompts of some systems to adapt to our selected tasks. The modified prompts are detailed in §B.1 of the Appendix. We use GPT-3.5 and GPT-4o as the backbone with a temperature of zero for main experiments (RQ1 and RQ2) while using GPT-3.5 for the remaining. Our findings are consistent with the GPT-4o results; see §A.1 in the appendix for details, as they are omitted here due to space constraints. We introduce one malicious agent at a time to avoid interference and facilitate essential analysis. Non-malicious agents remain unaware of the malicious agent’s presence, reflecting a realistic information-asymmetric scenario Zhou et al. (2024a).
4.2 RQ1: Impact of System Architectures
@@
AutoInject causes a larger performance drop than AutoTransform. While one might assume AutoTransform would have a greater negative impact on multi-agent collaboration due to its permanent modification of agents’ profiles into malicious ones, it is AutoInject that results in a more significant performance drop, although AutoInject introduces errors into a fixed and relatively small portion of messages. The reasons for this are two-fold: (1) Current LLMs have a weakness where they become less effective as the context lengthens, especially where conflict exists in instructions. For our malicious agents, they gradually lose track of the task to produce errors, prioritizing new instructions from other agents to correct errors in the message. (2) AutoInject consistently introduces errors, whereas AutoTransform does not always ensure error generation. Despite being transformed into malicious agents, they sometimes fail to generate errors due to constraints requiring errors to be stealthy.
(a) The three multi-agent system architectures.
(b) The four downstream tasks.
Figure 3: Performance drops of the six multi-agent systems on four selected downstream tasks.

### v3

# On the Resilience of LLM-Based Multi-Agent Collaboration with Faulty Agents

@@ Abstract
Multi-Agent System, Large Language Model, Resilience, Robustness
@@ 1 Introduction
Multi-agent collaboration has further boosted Large Language Models’ (LLMs) already impressive performance across various downstream tasks, including code generation (Liu et al., 2023; Lee et al., 2024), math problem solving (Lu et al., 2024; Liang et al., 2024), and text translation (Jiao et al., 2023; Wu et al., 2024a). Multi-agent systems achieve these improvements by decomposing complex tasks into smaller, specialized sub-tasks handled by expert agents (Chen et al., 2024b; Li et al., 2024).
Figure 1: We focus on the overall impact of faulty agents on the performance of diverse system structures across various tasks.
@@
However, the decentralized nature of multi-agent systems leaves them vulnerable to clumsy or malicious agents, which could undermine or destroy collaboration (Chen et al., 205). Consider a scenario where companies specializing in different areas produce expert agents, the lack of centralized control means that the multi-agent system may contain agents from various sources, some of which could be faulty. In a multi-agent coding system like Camel (Li et al., 2023), a faulty coding agent produces buggy code, causing severe errors or harmful outputs when executed by other agents.
@@
Recent studies (Zhang et al., 2024; Tian et al., 2023; Amayuelas et al., 2024; Ju et al., 2024; Yu et al., 2024) have increasingly focused on safety issues within multi-agent systems. However, these studies mainly investigate attacks on agents to induce toxicity in their outputs or misinformation spread among all agents. While they assess malicious agent behavior against safety benchmarks like AdvBench (Zou et al., 2023), they overlook the disruption of collaboration in solving general tasks and the impact of varying system structures. In this paper, we study the resilience of multi-agent collaboration against faulty agents, specifically the systems’ ability to recover from errors.
@@
First, we propose two approaches to simulating agents’ faulty behaviors on various tasks, namely AutoTransform and AutoInject. AutoTransform transforms a given agent’s profile into a faulty version that retains original functionalities while introducing stealthy errors. AutoInject is designed to directly and automatically inject errors into messages spread among agents. The two methods offer automate introduction of errors in multi-agent collaboration without requiring manual modifications.
@@
Then, we study the macro-level impact of faulty agents in different system structures and downstream tasks, particularly how their presence leads to an overall performance decline. We select six multi-agent collaboration systems that represent three classical human organizational structures: Linear (Hong et al., 2024; Dong et al., 2024), Flat (Li et al., 2023; Wang et al., 2024c), and Hierarchical (Chen et al., 2024b; Liang et al., 2024). We evaluate the performance of these systems across four tasks: code generation (Chen et al., 2021), math problem solving (Liang et al., 2024), translation (He et al., 2020), and text evaluation (Wang et al., 2024a), as shown in Fig. 1. Additionally, we analyze the impact of different error types (semantic or syntactic) and error rates on overall system resilience in code generation.
@@
Finally, we introduce two strategies for enhancing system resilience and recovering from faulty agents, each inspired from one of the proposed error-introducing methods. The “Challenger” method adds to each agent’s profile the ability to challenge received messages, mirroring AutoTransform which rewrites agents’ profiles to make them faulty. The “Inspector” is an extra agent who reviews and corrects messages, mirroring AutoInject which intercepts and injects errors into messages.
Our key findings include: (1) The Hierarchical structure exhibits the least performance drop at $9.2\%$, aligning with its prevalence in human organizational structures (Mihm et al., 2010). (2) Code Generation, as a relatively objective task, is most affected by malicious agents, experiencing a performance drop of $24.7\%$. (3) Manually introducing errors can sometimes improve the overall performance, especially on MAD (Liang et al., 2024). (4) Increasing the ratio of Faulty Messages and using Semantic Errors results in a greater performance drop than increasing the number of errors per message and using syntactic errors. (5) The combination of The Challenger and The Inspector enhances system resilience most for the two more vulnerable systems: Self-collab with a linear structure and Camel with a flat structure, recovering up to $96.4\%$ of performance lost caused by faulty agents. The contribution of this paper are as follows:
@@
  • We are the first to examine how different structures of multi-agent systems affect resilience when faulty agents exist and disrupt collaboration.
  • We design AutoTransform and AutoInject to automatically simulate agents’ faulty behaviors, and the Inspector and the Challenger to improve system resilience.
@@
  • We conduct extensive experiments involving six multi-agent systems across three system structures, applied to four common downstream tasks, offering detailed insights into designing resilient multi-agent systems.
@@ 2 Preliminaries
Collaboration: A Management Science Perspective
@@
Figure 2: Overview of our error-introducing methods. (a) Task information. (b) Multi-agent collaboration system without faulty agents. (c) AutoTransform modifies agent’s profile to turn it into faulty while preserving original functionalities. (d) AutoInject intercepts messages between agents and adds errors into the messages.
@@
In human collaboration, the capacity to handle internal errors and maintain overall operation without being affected by a single failure is usually referred to as “resilience” (Alliger et al., 2015; Boin & Van Eeten, 2013; Hartwig et al., 2020). LLM-based multi-agent collaboration faces robustness issues when clumsy or even malicious agents produce errors too stealthy to be found by other agents but can cause undesired consequences. Therefore, holding this same ability as human collaboration to recover from errors becomes critical.
@@ 3 Methodology: Introducing Errors
We offer two methods for introducing errors in multi-agent systems: AutoTransform converts agents into faulty ones that generate errors autonomously, while AutoInject directly introduces errors into messages. In this section, we first discuss the design of the autonomous transformation aproach in §3.1. Next, we introduce the method for directly injecting errors into messages within multi-agent collaboration in §3.2. These two methods are designed to be general-purpose, applicable to any agent profiles and downstream tasks. For presentation clarity, we use “message” to refer to intermediate outputs between agents, and “result” to denote the final output from the last agent.
3.1 AutoTransform: Faulty Agent Transformation
AutoTransform is an LLM-based approach that takes any agent’s profile as input and outputs a profile of a faulty agent performing the same functions but introducing stealthy errors. Drawing inspiration from how we manually convert an agent into malicious one, the design of AutoTransform follows three key steps: (1) To ensure applicability to any target agent and downstream tasks, AutoTransform first analyzes the input agent profile and extract the assigned task. This step helps to understand the task and identify potential ways to produce erroneous outputs. (2) Based on the task analysis, AutoTransform lists all possible methods to introduce errors, emphasizing the need for stealth to avoid detection by other agents. (3) AutoTransform then rewrites the agent’s profile with these error-injection methods, ensuring that the original functionalities of the agent remain unchanged. An example of using AutoTransform to modify an agent’s profile is shown in Fig. 2c. The complete prompt is provided in §C.3.
@@ 3.2 AutoInject: Direct Error Injection
While AutoTransform can conveniently generate malicious agents, it is hard to ensure these agents introduce a specific number and type of errors due to the inherent randomness of LLMs’ generation process. For example, “injecting syntax errors in 20% lines of the generated code” cannot be guaranteed by the faulty agents. However, precise error generation is crucial for analyzing the impact of various factors on system resilience. To address this, we introduce AutoInject, an approach that takes the outputs of other agents and intentionally injects specific errors. This approach allows for exact control over the proportion of erroneous messages, the specific errors within a message, and the types of errors introduced. We start by discussing two key factors in our study: error rate and error type.
@@
We examine two aspects of error injection in multi-agent collaboration systems: Macro Perspective: We control the ratio of erroneous messages produced by a faulty agent in all its messages, which is a practical way to obscure its incompetent identity while facilitating stealthy errors. We denote this probability that a message is intentionally flawed as $P_{m}$. Micro Perspective: We manage the degree of error within each faulty message. For instance, in code generation tasks, we can adjust the number of lines of erroneous code. The proportion of errors in a message is denoted by $P_{e}$.
@@
AutoInject requires inputs including task specifications, agent details, error rates ($P_{m}$ and $P_{e}$, defaulting to 1.0 and 0.2, respectively), and error type, which defaults to semantic errors. It then selects messages from the agent with a probability of $P_{m}$ and injects errors into $P_{e}$ of the total lines or sentences in the selected message. Errors are introduced automatically using LLMs, which receive the task introduction, error type, and the specific line or sentence to produce erroneous lines or sentences, replacing the originals. An example of using AutoInject to modify an agent’s output into erroneous is shown in Fig. 2d. Prompts for different tasks are detailed in §C.4.
@@ 4 Experiments
This section focuses on answering the following research questions: (1) Which of the three multi-agent system structures exhibits the highest resilience (§4.2)? (2) Do different downstream tasks vary in their resilience to errors (§4.3)? (3) How do varying error rates (both $P_{m}$ and $P_{e}$) impact system resilience (§4.4)? (4) How do the two types of errors influence system resilience (§4.5)?
(a) Backbone LLM: GPT-3.5.
(b) Backbone LLM: GPT-4o.
Figure 3: The performance of various system structures with the two error-introducing methods, with results averaged across all four tasks.
4.1 Experimental Settings
@@
We assess four tasks that evaluate general-purpose problem-solving abilities. All the evaluation metrics range from 00 to $100$ with higher values indicating better performance, allowing us to compute the overall performance by averaging scores across the four tasks.
  • Code Generation: HumanEval (Chen et al., 2021) contains 164 hand-written programming problems to assess LLMs’ ability to synthesize correct and functional Python code. Accuracy (Pass@1) is used for evaluation.
  • Math Problem Solving: CIAR (Liang et al., 2024) presents 50 questions with hidden traps to evaluate LLMs’ Counter-Intuitive Arithmetic Reasoning abilities, requiring multi-step reasoning. Accuracy is used for evaluation.
  • Translation: CommonMT (He et al., 2020) consists of paired sentences to test models’ handling of three types of commonsense reasoning, especially in ambiguous contexts. We randomly sampled 100 sentences from the most challenging type, Lexical, for our evaluation, using BLEURT-20 (Sellam et al., 2020; Pu et al., 2021) for evaluation, following the practice in Liang et al. (2024).
  • Text Evaluation: FairEval (Wang et al., 2024a) includes 80 human-annotated “win/tie/lose” labels comparing responses from ChatGPT and Vicuna-13B, aiming to determine if the model’s preferences align with human judgments. Accuracy is used for evaluation.
Type Name Tasks Num. Final Agent Faulty Agent
Linear MetaGPT All 5 Test Engineer Code Engineer
Self-collab Code 3 Tester Coder
[…]
SPP Code 2$\sim$5 AI Assistant Programmer
[…]
AgentVerse All 4 Critic Solver
Table 1: Details of the six multi-agent systems. “Num.” is the number of agents. “Final Agent” denotes the agent that output the final results.
@@
We use six multi-agent systems for the three types of structures mentioned in §2:
  • Linear: MetaGPT (Hong et al., 2024) uses Standard Operating Procedures (SOPs) to create an efficient workflow in a company of five agents. Self-collaboration (Dong et al., 2024) designs three roles, namely analyzers, coders, and testers, for code generation.
  • Flat: Camel (Li et al., 2023) presents a framework where a “User” agent iteratively refines outputs from an “Assistant” agent, applicable across various tasks. SPP (Wang et al., 2024c) uses Solo-Performance-Prompting to engage a single model into 2$\sim$5 personas for coding tasks.
  • Hierarchical: MAD (Liang et al., 2024) introduces a Multi-Agent Debate framework with two debaters and one judge to promote divergent thinking in LLMs. AgentVerse (Chen et al., 2024b) employs a dynamic recruitment process, selecting agents for multi-round collaboration as needed, using four agents in our selected tasks.
Not all systems are designed to support the four tasks studied in our paper. Therefore, we modified the prompts of some systems to adapt to our selected tasks. The modified prompts are detailed in §C.1. GPT-3.5 is consistently used for both AutoTransform and AutoInject to ensure a fair comparison. We use GPT-3.5 and GPT-4o as the backbone for these systems for main experiments (RQ1 and RQ2) while using GPT-3.5 for factor analysis. All LLMs are used with a temperature of zero. We introduce one faulty agent at a time to avoid interference and facilitate essential analysis, which is shown in Table 1. Normal agents remain unaware of the faulty agent’s presence, reflecting a realistic information-asymmetric scenario (Zhou et al., 2024a).
4.2 RQ1: Impact of System Structures
@@
AutoInject causes a larger performance drop than AutoTransform on GPT-3.5, but a lower performance drop using GPT-4o. While one might assume AutoTransform would have a greater negative impact on multi-agent collaboration due to its permanent modification of agents’ profiles into faulty ones, it is AutoInject that results in a larger performance drop using GPT-3.5. The reasons for this are two-fold: (1) Current LLMs have a weakness where they become less effective as the context lengthens, especially where conflict exists in instructions. For our faulty agents, they gradually lose track of the task to produce errors, prioritizing new instructions from other agents to correct errors in the message. (2) AutoInject consistently introduces errors, whereas AutoTransform does not always ensure error generation. Despite being transformed into faulty agents, they sometimes fail to generate errors due to constraints requiring errors to be stealthy. These issues are mitigated as the capabilities of LLMs advance. With GPT-4o as the backbone, the faulty agents generated by AutoTransform demonstrate a strong capacity for instruction following, resulting in stealth errors that lead to a more significant performance decline compared to AutoInject.
(a) Backbone LLM: GPT-3.5.
(b) Backbone LLM: GPT-4o.
Figure 4: The performance of various tasks with the two error-introducing methods, with results averaged across all six multi-agent systems (three system structures).